OpenStudy (anonymous):

find 4 1/5 x 7 2/3 a)28 2/15 b)28 3/10 c)32 1/5 d)32 3/5

OpenStudy (johnweldon1993):

We should change them into improper fractions instead of mixed numbers \[4 \frac{ 1 }{ 5 } = \frac{ 21 }{ 5 }\] and \[7 \frac{ 2 }{ 3 } = \frac{ 23 }{ 3 }\] Now we have \[\frac{ 21 }{ 5 } \times \frac{ 23 }{ 3 }\] So this comes out to \[\frac{ 483 }{ 15 }\] Can you turn this back into a mixed fraction?

OpenStudy (johnweldon1993):

To turn it back into a mixed fraction...you figure out how many times 15 goes into 483 evenly..... this comes out to 32 because 32 times 15 = 480 so now we see what we have leftover....483 - 480 = 3 And we put that over the original denominator so this is \[32 \frac{ 3 }{ 15 }\] But this can be simplified into \[32 \frac{ 1 }{ 5 }\]